How many quarts of a 10% antifreeze solution must be mixed with 16 quarts of a 40% antifreeze solution to make a 30% solution.
Answer by mananth(16946)

------ percent ---------------- quantity
Antifreeze I 40 ---------------- 16 quarts
Antifreeze II 10 ---------------- x quarts
Total 30 ---------------- 16 + x quarts

40*16+10 *x=30(16 +x)
640+10x= 480+30x
10x-30x= 480-640
-20x=-160
/-20
x=8 quarts 10% Antifreeze II